OpenStudy (anonymous):

The triangular sail on Richard's boat needs to be replaced. Find the amount of fabric needed for a new sail that has a base of 16 meters and a height of 9 meters.

OpenStudy (anonymous):

The triangular sail on Richard's boat needs to be replaced. Find the amount of fabric needed for a new sail that has a base of 16 meters and a height of 9 meters. NOTE :TRIANGULAR SAIL NEEDS TO BE REPLACED. complete area it is a triangle shape use area of a triangle formula there is the height and base given this is all you need to solve

OpenStudy (anonymous):

what is the triangle formula?

OpenStudy (anonymous):

the formula is A = 1/2bh Area = 1/2 * base of triangle * height of triangle